Show whether the first three vectors are linearly independent

V1= 1, 1, -2, -2

V2= 2, -3, 0, 2

V3= -2, 0, 2, 2

V4= 3, -3 -2, 2
1
Expert's answer
2020-11-12T17:40:44-0500

"\begin{cases}
 x_1+2x_2-2x_3=0, (1)\\ 
 x_1-3x_2=0, (2)\\
-2x_1+2x_3=0, (3)\\
-2x_1+2x_2+2x_3=0. (4)
\end{cases}"

From (2): "x_2=x_1\/3", from (3): "x_3=x_1," to (1):

"x_1+2x_1\/3-x_1=0,"

"x_1=0,"

(2): "x_2=0,"

(3): "x_3=0."

Only "x_1=x_2=x_3=0" is the solution.

So, by the defenition, vectors are linearly independent.